OpenStudy (anonymous):

Solve for x: (2x-7)(x+4)=0

OpenStudy (anonymous):

2x-7=0 x+4=0 2x=7 x=-4 x=3.5 or x=-4

OpenStudy (anonymous):

Either 2x-7=0 or x+4=0 Now point 1: 2x-7=0 which implies 2x=7 which gives x=7/2 Now point 2: x+4=0 which gives x=-4

OpenStudy (anonymous):

i for some reason got 7/2 -4